# InterFormer: Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v3

@@ Abstract.
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an ad, is a fundamental task in recommender systems. The emergence of heterogeneous information, such as user profile and behavior sequences, depicts user interests from different aspects. A mutually beneficial integration of heterogeneous information is the cornerstone towards the success of CTR prediction. However, most of the existing methods suffer from two fundamental limitations, including (1) insufficient inter-mode interaction due to the unidirectional information flow between modes, and (2) aggressive information aggregation caused by early summarization, resulting in excessive information loss. To address the above limitations, we propose a novel module named InterFormer to learn heterogeneous information interaction in an interleaving style. To achieve better interaction learning, InterFormer enables bidirectional information flow for mutually beneficial learning across different modes. To avoid aggressive information aggregation, we retain complete information in each data mode and use a separate Cross Arch for effective information selection and summarization. The proposed InterFormer has been deployed across multiple platforms at Meta Ads, achieving 0.15% performance gain and 24% QPS gain compared to prior state-of-the-art models and yielding sizable online impact.
@@ 1. Introduction
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an item, is the fundamental task for various applications such as online advertising and recommender systems (Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences company revenue and user experience, drawing extensive attention from academia and industry (Zhou et al., 2018, 2019). For example, in online ad bidding, CTR prediction helps advertisers optimize the bids and target the most receptive audiences. In content recommendation, it enables platforms to suggest relevant content to users.
@@ 2. Related Works
In this section, we briefly review the related works on feature interaction, including non-sequential and sequential methods.
@@ 2.2. Sequential Methods
With the recent emergence of sequential information, e.g., user interaction history, in recommender systems, extensive efforts have been made to capture the evolving user interest. A key challenge behind sequential methods is to combine the sequential information with non-sequential information in a mutually beneficial manner. Markov models (Shani et al., 2005; He and McAuley, 2016a; Yang et al., 2020) consider sequential data as a stochastic process over discrete random variables, but the oversimplified Markovian property limits the model capability in capturing long-term sequential dependencies (Quadrana et al., 2018). To model the long-term dependencies, RNN and Attention mechanism are employed as the backbone module for many sequential methods. For example, various attention-based networks are designed (Zhou et al., 2019; Lyu et al., 2020), and Transformer (Devlin et al., 2018; Vaswani et al., 2017) is adopted for sequential modeling (Sun et al., 2019; Chen et al., 2019). Besides, to model multi-faceted user interests, (Xiao et al., 2020; Han et al., 2024) propose to capture multiple user interests from multi-behavior sequences. More recently, TransAct (Xia et al., 2023) adopt a hybrid ranking model to combine real-time user actions for immediate preference and batch user representations for long-term interests. LiRank (Borisyuk et al., 2024) improves CTR prediction at LinkedIn by ensembling multiple interaction modules, accelerated by quantization and vocabulary compression. CARL (Chen et al., 2024) achieves fast inference on large-scale recommendation in Kuaishou by utilizing cached results, monitored by a reinforcement learning-based framework. While most of the existing sequential methods leverage non-sequential information for personalized sequence modeling, sequential information is rarely explored for non-sequence learning. We believe such unidirectional design limits the expressiveness of the learned embeddings, and a bidirectional information flow between different modes is the key towards better heterogeneous interaction learning.
@@ 3. Preliminaries
[…]
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
(8) $\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)=f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)\mathbf{S}^{(l)},\vspace{-5pt}$
@@
(9) $\mathbf{S}^{(l+1)}=\text{MHA}^{(l)}\left(\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)\right).\vspace{-5pt}$
@@ 4.4. Cross Arch: Towards Effective Information Selection and Summarization
where $\text{where }\text{Gating}(\mathbf{X})=\sigma\left(\mathbf{X}\odot\text{MLP}(\mathbf{X})\right).$
@@
(11) $\mathbf{S}^{(l)}_{\text{sum}}=\text{Gating}\left(\left[\mathbf{S}^{(l)}_{\text{CLS}}\|\mathbf{S}^{(l)}_{\text{PMA}}\|\mathbf{S}^{(l)}_{\text{recent}}\right]\right).$
@@ 5.1. Experiment Setup
  • LogLoss (cross-entropy loss) measures the distance between the prediction $\hat{y}$ and the label $y$, and is computed as $L(y,\hat{y})\!=\!-\left(y\log\left(\hat{y}\right)+(1\!-\!y)\log\left(1\!-\!\hat{y}\right)\right)$. Lower the better.
@@ A.1. Algorithm
9:  Compute CTR prediction by $\hat{y}=\text{MLP}\left(\left[\mathbf{X}^{(L)}_{\text{sum}}\|\mathbf{S}^{(L)}_{\text{sum}}\right]\right)$;
@@ A.2.1. Linear Compressed Embedding
Given numerous non-sequence features, it is essential to compress embeddings to a manageable size. In general, given $N$ $d$-dimensional features $\mathbf{X}\in\mathbb{R}^{d\times N}$, LCE is a linear transformation on the sample dimension $\mathbf{W}\in\mathbb{R}^{N\times M}$, such that the linear transformation $\mathbf{XW}\in\mathbb{R}^{d\times M}$ serves as the compressed embedding with $M$ features. Together with self-gating modules, LCE selectively compress and denoise numerous embeddings, benefiting the model in both efficiency and effectiveness.

### v4

@@ Abstract.
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an ad, is a fundamental task in recommender systems. The emergence of heterogeneous information, such as user profile and behavior sequences, depicts user interests from different aspects. A mutually beneficial integration of heterogeneous information is the cornerstone towards the success of CTR prediction. However, most of the existing methods suffer from two fundamental limitations, including (1) insufficient inter-mode interaction due to the unidirectional information flow between modes, and (2) aggressive information aggregation caused by early summarization, resulting in excessive information loss. To address these limitations, we propose a novel module named InterFormer to learn heterogeneous information interaction in an interleaving style. To achieve better interaction learning, InterFormer enables bidirectional information flow for mutually beneficial learning across different modes. To avoid aggressive information aggregation, we retain complete information in each data mode and use a separate Cross Arch for effective information selection and summarization. InterFormer has been deployed across multiple platforms at Meta Ads, achieving 0.15% performance gain and 24% QPS gain compared to prior state-of-the-art models and yielding sizable online impact.
@@ 1. Introduction
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an item, is the fundamental task for various applications such as online advertising and recommender systems (Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences company revenue and user experience, drawing extensive attention from academia and industry (Zhou et al., 2018, 2019; Liu et al., 2024b; Liang et al., 2025; Yoo et al., 2024; Liu et al., 2023, 2024a, 2025). For example, in ad bidding, CTR prediction helps advertisers optimize the bids and target the most receptive audiences. In content recommendation, it enables platforms to suggest relevant content to users.
@@ 2. Related Works
In the era of big data and AI (Yan et al., 2021a, b, 2023a, 2023b, 2022, 2024b, 2024c, 2024a, [n. d.]; Xu et al., [n. d.]; Yu et al., 2025b, a; Zeng et al., [n. d.], 2025, 2023a, 2023b, 2024; Bao et al., 2024; Xu et al., 2024; Lin et al., 2024, 2025b, 2025a), recommender systems have drawn significant attention, exerting profound influence on company revenue and user experience (Wang et al., 2018, 2023b, 2017, 2021b; Li et al., 2022, 2024; Lee et al., 2019; Jing et al., 2024, 2022; Yang et al., 2024; Wang et al., 2023a). In this section, we review the related works on CTR prediction, including non-sequential and sequential methods.
@@ 2.2. Sequential Methods
With the recent emergence of sequential information, e.g., user interaction history, in recommender systems, extensive efforts have been made to capture the evolving user interest. A key challenge behind sequential methods is to combine the sequential information with non-sequential information in a mutually beneficial manner. Markov models (Shani et al., 2005; He and McAuley, 2016a; Yang et al., 2020) consider sequential data as a stochastic process over discrete random variables, but the oversimplified Markovian property limits the model capability in capturing long-term sequential dependencies (Quadrana et al., 2018). To model the long-term dependencies, RNN and Attention mechanism are employed as the backbone module for many sequential methods. For example, various attention-based networks are designed (Zhou et al., 2019; Lyu et al., 2020), and Transformer (Devlin et al., 2018; Vaswani et al., 2017; Lin et al., 2025c) is adopted for sequential modeling (Sun et al., 2019; Chen et al., 2019). Besides, to model multi-faceted user interests, (Xiao et al., 2020; Han et al., 2024) propose to capture multiple user interests from multi-behavior sequences. More recently, TransAct (Xia et al., 2023) adopt a hybrid ranking model to combine real-time user actions for immediate preference and batch user representations for long-term interests. LiRank (Borisyuk et al., 2024) improves CTR prediction at LinkedIn by ensembling multiple interaction modules, accelerated by quantization and vocabulary compression. CARL (Chen et al., 2024) achieves fast inference on large-scale recommendation in Kuaishou by utilizing cached results, monitored by a reinforcement learning-based framework. While most of the existing sequential methods leverage non-sequential information for personalized sequence modeling, sequential information is rarely explored for non-sequence learning. We believe such unidirectional design limits the expressiveness of the learned embeddings, and a bidirectional information flow between different modes is the key towards better heterogeneous interaction learning.
@@ 3. Preliminaries
$\langle\cdot,\cdot\rangle$ inner product
$\cdot\|\cdot$] horizontal concatenation of vectors
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
(8) $\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)=f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)\mathbf{S}^{(l)},\vskip-5.0pt$
@@
(9) $\mathbf{S}^{(l+1)}=\text{MHA}^{(l)}\left(\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)\right).\vskip-5.0pt$
@@ 4.4. Cross Arch: Towards Effective Information Selection and Summarization
$\text{where }\text{Gating}(\mathbf{X})=\sigma\left(\mathbf{X}\odot\text{MLP}(\mathbf{X})\right).$
@@
[…]
@@ 5.1. Experiment Setup
[…]
@@ A.1. Algorithm
[…]
@@ A.2.1. Linear Compressed Embedding
Given numerous non-sequence features, it is essential to compress embeddings to a manageable size. In general, given NN $d$-dimensional features $\mathbf{X}\in\mathbb{R}^{d\times N}$, LCE is a linear transformation on the sample dimension $\mathbf{W}\in\mathbb{R}^{N\times M}$, such that the linear transformation $\mathbf{XW}\in\mathbb{R}^{d\times M}$ serves as the compressed embedding with $M$ features. Together with self-gating modules, LCE selectively compress and denoise numerous embeddings, benefiting the model in both efficiency and effectiveness.